OpenStudy (ineedatutor):

PLEASE help! I don't want the answers, I just need to know how to do this! 2. Ramone and his wife had $72,380.90 in combined gross income in 2010. They had $8,295.00 in various adjustments and deductions. If they are filing jointly, how much income tax do they have to pay for the year?

OpenStudy (ineedatutor):

Here's the table

OpenStudy (sshayer):

Taxable income=gross income-deductions then show me your work and i will guide further.

OpenStudy (ineedatutor):

@sshayer would that be $72,380.90-$8,295.00?

OpenStudy (sshayer):

correct.

OpenStudy (sshayer):

calculate taxable income.

jimthompson5910 (jim_thompson5910):

Let x = taxable income = difference of gross income and deductions T = income tax paid = amount sent to the government Use the "filing jointly" chart on the right side. You'll use the second row of that chart since subtracting `$72,380.90-$8,295.00` leads to a value between 16750 and 68000. That second row turns into this equation T = 1675 + (x-16750) where \(\Large 16750 < x \le 68000\)

jimthompson5910 (jim_thompson5910):

sorry, I meant to write this equation T = 1675 + 0.15*(x-16750)

OpenStudy (ineedatutor):

I think I'm understanding now.

OpenStudy (ineedatutor):

@jim_thompson5910 would I use 64085.9 for X?

jimthompson5910 (jim_thompson5910):

`would I use 64085.9 for X?` yes @IneedaTutor